# Jan Schiefelbein-Lach4, Oliver Pohl5, Andreas Ulbig2, and Martin Grohe1

### v1

@@ Abstract
In the course of the energy transition, the expansion of generation and consumption will change, and many of these technologies, such as PV systems, electric cars and heat pumps, will influence the power flow, especially in the distribution grids. Scalable methods that can make decisions for each grid connection are needed to enable congestion-free grid operation in the distribution grids. This paper presents a novel end-to-end approach to resolving congestion in distribution grids with deep reinforcement learning. Our architecture learns to curtail power and set appropriate reactive power to determine a non-congested and, thus, feasible grid state. State-of-the-art methods such as the optimal power flow (OPF) demand high computational costs and detailed measurements of every bus in a grid. In contrast, the presented method enables decisions under sparse information with just some buses observable in the grid. Distribution grids are generally not yet fully digitized and observable, so this method can be used for decision-making on the majority of low-voltage grids. On a real low-voltage grid the approach resolves 100% of violations in the voltage band and 98.8% of asset overloads. The results show that decisions can also be made on real grids that guarantee sufficient quality for congestion-free grid operation.
@@
Smart grids, Edge Computing, State Estimation, Optimal Power Flow, Deep Reinforcement Learning
**footnotetext: These authors contributed equally to this workfootnotetext: Funded by the European Union (ERC, SymSim, 101054974). Views and opinions expressed are, however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ I Introduction
As part of the decarbonization of energy systems, new loads and generators are being connected to the distribution grids. The majority of electricity generation plants, such as photovoltaic and wind power plants, are being built decentrally. Furthermore, heat pumps and electromobility are replacing technologies that were previously based on fossil fuels. This changes and increases the power flow in the electricity grids, especially in the distribution grids.
To cope with this issue, customers’ power demand or generation can be curtailed in cases of congestion. Independently of regulatory constraints, the need to decide which appliances connected to the grid must be curtailed to resolve the congestion exists. Solving an Optimal Power Flow (OPF) is the state-of-the-art approach to this question. However, to solve an OPF, measurements at every bus in the grid are needed. Often, such measurements are not available. A standard way of coping with missing measurements is to use state estimation techniques to predict the grid state on the basis of a few measurements and then apply optimal power flow to the predicted grid state. Current developments show that predicting grid states by using state estimation is a measure for distribution system operators to obtain complete information about their grid states. Unfortunately, state estimation is computationally expensive and introduces some inaccuracy.
Our approach, on the other hand, shows a novel method for obtaining curtailment decisions in power grids in an end-to-end fashion. Based on a few measurements, we utilize a model to predict curtailment decisions with high accuracy, which is deployable on low-cost edge computing devices. Our method uses the possibilities of reinforcement learning with partial observability by iteratively learning the curtailment decisions to obtain non-critical grid states. Reinforcement learning under partial observability finds normally application, e.g. in robotics [1], and video games [2].
@@ I-A Related work
Recent studies have demonstrated an increasing interest in utilizing Machine Learning (ML) techniques to estimate the solution of AC-OPF. This is achieved either by supervised learning, through self-supervised approaches or using Reinforcement Learning (RL) paradigms. Some of the early works relied on constrained supervised training. The work [3] incorporated the Karush-Kuhn-Tacker (KKT) conditions as penalty terms during training, while [4] combines a Multi-Layer Perceptron (MLP) with the Lagrangian Dual method. The effort in [5] leverages graph learning and implements a local and a global Graph Neural Network (GNN) to solve the OPF problem. Although supervised approaches may yield promising results, they require the generation of pre-solved instances for training, which may be time-consuming and resource-intensive. Other approaches mitigate this drawback by relying on self-supervised training. A follow-up work to the method in [5] is presented in [6]. The key difference from the previous work is that the new method is trained using an unsupervised algorithm and a barrier method to satisfy the constraints. The work DC3 presented in [7] implements an MLP, where a completion step based on Newton’s Method is performed, and feasibility of the inequality constraints is ensured using a gradient-based correction step. On the other hand, other works exploit the self-learning ability of RL-agents and employ similar techniques for constrained learning to direct the RL agent in making physics-informed decisions. [8] and [9] rely on adding system specific Lagrangian parameters to the reward function, whereas the works in [10] and in [11] use the safety convex layer and holomorphic embedding based layer respectively to ensure feasibility.
@@ I-B Main Contribution
This paper introduces a novel approach that leverages deep reinforcement learning to determine the level of curtailment for controllable buses, relying on a partially observable grid state. The method utilizes available grid measurements as inputs to a reinforcement learning agent, which directly determines the curtailment level for each bus. The agent undergoes training, following the typical reinforcement learning paradigm, within an environment that simulates the grid under the current supply task using power flow computations. To validate the proposed method, real low-voltage grid data from Schleswig-Holstein Netz AG and synthetic data representing realistic supply tasks are employed for training and validation. Results showcase the model’s efficacy in resolving nearly all instances of violations of physical constraints. Instances where violations persist are correctly identified by the model, however with insufficient curtailment
II Background
@@
Optimal Power Flow (OPF) combines the economic dispatch of generation and load and the physical constraints through the power flow formulations. Optimal power flow is a nonlinear and non-convex, NP-complete, constrained optimization problem on power grids, commonly used in areas such as grid planning, power markets, and active grid operation and is a backbone in many calculations used to operate modern power systems [12, 13]. For that reason, there exist different types and modifications of the standard AC OPF problem, e.g., DC OPF, security-constrained OPF and others [14]. The following formulation describes the Standard OPF problem. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a grid with buses $\mathcal{V}$ and lines $\mathcal{E}$.
@@
III Method
The architecture needs a fixed grid topology with multiple supply tasks, which model different grid states including violations of lower or upper voltage bands, as well as line overloads. The architecture shall learn to resolve these violations by curtailing active and reactive power at certain buses with flexibility, which are explicitly listed as controllable. However, in reality, not all of the buses have measurement equipment installed. Therefore, the architecture only gets measurements of observable buses. We assume that every controllable bus is also observable. At the observable buses, measurements of $P,Q$ and $V$ are taken. Additionally, we supply valid ranges $[P_{i}^{\text{min}},P_{i}^{\text{max}}]$ and $[Q_{i}^{\text{min}},Q_{i}^{\text{max}}]$ for each controllable bus $i$. The architecture shall learn the optimal setpoint for $P$ and $Q$ at the controllable buses based on the restricted measurements (or observation).
III-A Data Generation
To create datasets, it is necessary to map different grid operating states on which decisions have to be made. The dataset should encompass both non-critical and critical operational states to enable effective learning during training. By including a variety of scenarios, ranging from normal to critical states, the model can discern patterns and distinctions between the two. This ensures that the detection of critical states can be informed by the characteristics of non-critical states, facilitating a comprehensive learning of grid dynamics. The second aspect is the mapping of curtailment decisions in the grid so that flexible assets in the grid are curtailed in such a way that a non-critical grid state can be created from a critical grid state by the model’s decision.
Since, in many cases, grid models do not contain detailed information on the supply task, the first step is to assign an annual time series to the grid model. These are based on the technologies installed at the nodes, such as photovoltaics and household loads [15, 16]. The controllable installed technologies are assigned cost coefficients $c$ (cf. Eq. 1a) so that their optimal use is incited. In addition, observable and non-observable nodes are defined in the grid, which should represent nodes in reality with measurement technology.
Based on the modelled supply task on the grid, both a power flow and an optimal power flow on the grid are calculated. When calculating the optimal power flow, the grid is assumed to be completely observable. For comparison, a state estimation is also calculated in combination with an optimal power flow in order to take account of the incomplete observability and to obtain a reference for the model to be trained here.
@@ III-A Data Generation
The calculation is carried out for at least 35.040 quarter-hour values in order to represent an annual simulation. In the course of development, it has been shown that further augmentation of the reference time series on the basis of the annual simulation is expedient, in particular, to increase the number of critical grid states in the dataset in order to achieve a higher accuracy of the trained model. An example grid is presented in Fig. 1.
Figure 1: Low Voltage grid with 7% observability (provided by Schleswig Holstein Netz GmbH)
@@
The dataset, therefore, consists of tuples of unsolved and solved optimal power flow as well as an extension by a state estimation and optimal power flow on reduced observability.
III-B Environment
@@
The environment can be interpreted as a simulation of the physical world. It has a fixed grid topology, which gets updated with the corresponding supply task in every training step. The agent can interact with the environment by making an observation, i.e. collecting measurements at all observable buses. Based on this observation, the agent computes an action by choosing $P$ and $Q$ setpoints, which are applied to the grid. A power flow is computed with a Newton-Raphson procedure to simulate the effect of the agent’s decision. Based on the overall resulting state of the environment (which includes measurements at buses not observable by the agent), a reward is given to the agent.
III-C Reward
@@
Let $n$ be the number of buses, $k$ the number of controllable generators and $\ell$ be the number of lines. Let $\mathbf{V}$ be the vector of voltages of the buses in the current state, and denote by $\mathbf{V}^{\text{min}}$ and $\mathbf{V}^{\text{max}}$ the vector of lower and upper limits for the voltage band. We define the voltage loss $\mathcal{L}_{V}\coloneqq\max_{i\in\{1,\ldots,n\}}|\mathbf{V}-\text{clamp}(\mathbf{V},\mathbf{V}^{\text{min}},\mathbf{V}^{\text{max}})|_{i}$ as the maximum deviation from the voltage band. Formally, for vectors $\mathbf{x},\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ we define $\text{clamp}(\mathbf{x},\mathbf{a},\mathbf{b})\coloneqq\max(\mathbf{a},\min(\mathbf{x},\mathbf{b}))$ where $\min$ and $\max$ are applied pointwise. Similarly, for the vector of relative loads $\mathbf{I}$ we let $\mathcal{L}_{I}\coloneqq\max_{i\in\{1,\ldots,\ell\}}\max(\mathbf{I}-\mathbf{1},\mathbf{0})_{i}$ be the maximum relative line load violation. Further, we define the curtailment cost as $\mathcal{C}_{P}\coloneqq\frac{1}{|\mathbf{C}|}\sum_{i=1}^{|\mathbf{C}|}\mathbf{C}_{i}$ where $\mathbf{C}_{i}$ is the amount of power curtailed at the $i$-th controllable generator. The agent obtains a reward of -10 if the power flow (computed after applying the agent’s actions) does not converge. Otherwise, the agent obtains
@@
III-D Architecture
The agent that determines the $P$ and $Q$ setpoints is a neural network obtained from an Actor-Critic model, which we train using Deep Deterministic Policy Gradient (DDPG) [17]. Both Actor and Critic are simple multi-layer perceptrons with two hidden layers, each of width 512.
We have implemented the model in PyTorch [18] with TorchRL [19]. For the environment, we re-implemented pypower’s [20] Newton-Raphson in PyTorch for more efficient computation, as we can compute the power flow on the GPU. This has two advantages, the power flow computes faster, second we avoid copying data between GPU and main memory which has a tremendous overhead.
Figure 2: Relative P Curtailment against the physical limits. Each point in each box corresponds to one supply task. In the column baseline is the case where every flexible bus is on its maximum set point, the columns se+opf and rl contain each the solutions of state estimation and OPF and our model. The colour of each point corresponds to the sum of the flexibility of all busses in that supply task.
III-E Training Procedure
Reinforcement learning is often used when decisions from the past influence the future, e.g. balancing a pendulum facing upwards. In our setting, this is not clearly the case because the power flow in the grid can change heavily when a car is plugged into a charger, for example. So, from one time step to another, a decision that was made before could be totally wrong because customer behaviour resolved the congestion. Because of this, even though we are able to generate time-series data, the agent only sees the state at one timestep, and its actions depend on this one. For stability purposes, we hence shuffle the timesteps for training.
We train the agent’s policy based on experiences collected from the agent’s interactions with the environment. Since DDPG is an off-policy algorithm inherently deterministic, we add exploration noise (generated by an Ornstein-Uhlenbeck Process [21] with $\vartheta=0.355$ and $\sigma=0.6$) to the output of the Actor. Experiences are collected in a Replay Buffer of size 200,000 from which we sample 100 individual experiences uniformly at random at a time, which are then used to train the Actor and Critic networks.
IV Experiments
We conduct our experiments on a real low voltage grid of Schleswig-Holstein Netz AG (depicted in Figure 1), to which we apply our data generation method described in Section III-A. In cases where no flexibility is available on neither $P$ nor $Q$ at any bus, we dropped the supply task from the dataset since there is nothing to learn as $\mathbf{P}^{\text{min}}=\mathbf{P}^{\text{max}}$ and $\mathbf{Q}^{\text{min}}=\mathbf{Q}^{\text{max}}$. After this clean-up step, 20,453 supply tasks are still available. The distribution of violations is depicted in Table I.
TABLE I: Distribution of violations in test dataset (size: 20,453) and the ratio of resolved violations. Note that violations in either category are non-exclusive.

### v2

@@ Abstract
In the course of the energy transition, the expansion of generation and consumption will change, and many of these technologies, such as PV systems, electric cars and heat pumps, will influence the power flow, especially in the distribution grids. Scalable methods that can make decisions for each grid connection are needed to enable congestion-free grid operation in the distribution grids. This paper presents a novel end-to-end approach to resolving congestion in distribution grids with deep reinforcement learning (RL). Our RL-agent learns to curtail power and set appropriate reactive power to determine a non-congested and, thus, feasible grid state. State-of-the-art methods, such as the optimal power flow (OPF), require high computational costs and detailed measurements of every bus in a grid. In contrast, the presented method enables decisions under sparse information with just some buses observable in the grid. Distribution grids are generally not yet fully digitized and observable, so this method can be used for edge decision-making on the majority of low-voltage grids. On a real low-voltage grid, the approach resolves 100% of violations in the voltage band and 98.8% of asset overloads. The results show that decisions can also be made on real grids that guarantee sufficient quality for congestion-free grid operation.
@@
Curative Curtailment, Edge Computing, State Estimation, Optimal Power Flow, Deep Reinforcement Learning
publicationid: pubid: 979-8-3503-9042-1/ 24/\$31.00 ©2024 IEEE **footnotetext: These authors contributed equally to this workfootnotetext: Funded by the European Union (ERC, SymSim, 101054974). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ I Introduction
As part of the decarbonization of energy systems, new loads and generators are being connected to the distribution grids. The majority of electricity generation plants, such as photovoltaic are being built decentrally. Furthermore, heat pumps and electromobility are replacing technologies that were previously based on fossil fuels. This changes and increases the power flow in the electricity grids, especially in the distribution grids. To cope with this issue, the power demand or generation can be curtailed in cases of congestion to maintain the grids physical limits. Independently of regulatory constraints, the need to decide which appliances connected to the grid must be curtailed to resolve the congestion exists. Solving an Optimal Power Flow (OPF) is the state-of-the-art approach to this question. However, to solve an OPF, measurements at every bus in the grid are needed. Often, such measurements are not available. A standard way of coping with missing measurements is to use state estimation techniques to predict the grid state on the basis of a few measurements and then apply optimal power flow to the predicted grid state. Current developments show that predicting grid states by using state estimation is a measure for distribution system operators to obtain complete information about their grid states. Unfortunately, state estimation is computationally expensive and introduces some inaccuracy.
Our approach shows a novel method for obtaining curative curtailment decisions in power grids in an end-to-end fashion. Based on a few measurements, we utilize a model to predict curtailment decisions with high accuracy, which is deployable on low-cost edge computing devices. Our method uses the possibilities of reinforcement learning with partial observability by iteratively learning the curtailment decisions to obtain non-critical grid states. Reinforcement learning under partial observability finds normally application, e.g. in robotics [1].
@@ I-A Related work
Recent studies have demonstrated an increasing interest in utilizing Machine Learning (ML) techniques to estimate the solution of AC-OPF. This is achieved either by supervised learning, through self-supervised approaches or using Reinforcement Learning (RL) paradigms. Some of the early works relied on constrained supervised training. The work [2] incorporated the Karush-Kuhn-Tacker (KKT) conditions as penalty terms during training, while [3] combines a Multi-Layer Perceptron (MLP) with the Lagrangian Dual method. The effort in [4] leverages graph learning and implements a local and a global Graph Neural Network (GNN) to solve the OPF problem. Although supervised approaches may yield promising results, they require the generation of pre-solved instances for training, which may be time-consuming and resource-intensive. Other approaches mitigate this drawback by relying on self-supervised training. The work DC3 presented in [5] implements an MLP, where a completion step based on Newton’s Method is performed, and feasibility of the inequality constraints is ensured using a gradient-based correction step. On the other hand, other works exploit the self-learning ability of RL-agents and employ similar techniques for constrained learning to direct the RL agent in making physics-informed decisions. [6] and [7] rely on adding system-specific Lagrangian parameters to the reward function, whereas the works in [8] and in [9] use the safety convex layer and holomorphic embedding based layer respectively to ensure feasibility.
@@ I-B Contribution
This paper introduces a novel approach that leverages deep reinforcement learning to determine the level of curtailment for controllable buses, relying on a partially observable grid state. The method utilizes available grid measurements as inputs to a reinforcement learning agent, which directly determines the curtailment level for each bus. The agent undergoes training, following the typical reinforcement learning paradigm, within an environment that simulates the grid under the current supply task using power flow computations. To validate the proposed method, real low-voltage grid data from Schleswig-Holstein Netz AG and synthetic data representing realistic supply tasks are employed for training and validation. Results showcase the model’s efficacy in resolving nearly all instances of violations of physical constraints. Instances where violations persist are correctly identified by the model, however, with insufficient curtailment.
II Method
This section outlines the methodologies used to develop and train a model for grid curtailment with partial measurement availability. The model is capable of handling various operational scenarios ranging from non-critical to critical grid states. Central to our approach is the generation of a realistic dataset presented in II-A consisting of a wide spectrum of grid states and the application of a reinforcement learning approach presented in II-B to determine the curtailed power in critical situations.
The primary objective is to maintain the grid within permissible grid states by dynamically adjusting power outputs at controllable nodes within the grid. This involves both the direct control of power generation and the curtailment of power usage to prevent or resolve potential violations such as asset overloads and voltage band deviations. To achieve this, our method relies on a detailed environment that models the grid’s physical and operational constraints and employs an actor-critic architecture to optimize the decision-making process in an end-to-end concept.
II-A Data Generation
To create datasets, it is necessary to map different grid operating states on which decisions have to be made. The dataset should consist of both non-critical and critical operational states to enable effective learning during training. By including a variety of scenarios, ranging from normal to critical states, the model can determine patterns and distinctions between the two. This ensures that the detection of critical states can be informed by the characteristics of non-critical states, facilitating a comprehensive learning of grid dynamics. The second aspect is the mapping of curtailment decisions in the grid so that flexible assets in the grid are curtailed in such a way that a non-critical grid state can be created from a critical grid state by the model’s decision.
II-A1 Grid State Modelling
Since, in many cases, grid models do not contain detailed information on the supply task, the first step is to assign an annual time series to the grid model. These are based on the technologies installed at the nodes, such as photovoltaics and household loads [10, 11]. The controllable installed technologies are assigned realistic cost coefficients $c$ (cf. Eq. 1a) so that their optimal use is incited. In addition, observable and non-observable nodes are defined in the grid, which should represent nodes in reality with measurement technology. Based on the modelled supply task on the grid, an optimal power flow (cf. II-A2) on the grid is calculated.
II-A2 Optimal Power Flow
@@
The OPF used in this method combines the economic dispatch of generation and load and the physical constraints through the power flow formulations. Optimal power flow is a nonlinear and non-convex, NP-complete, constrained optimization problem on power grids, commonly used in areas such as grid planning, power markets, and active grid operation and is a backbone in many calculations used to operate modern power systems [12, 13]. For that reason, there exist different types and modifications of the standard AC OPF problem, e.g., DC OPF, security-constrained OPF, and others [14]. The following formulation describes the Standard OPF problem. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a grid with buses $\mathcal{V}$ and lines $\mathcal{E}$.
@@
II-A3 Training Data Set
@@
The calculation is carried out for at least 35.040 quarter-hour values in order to represent an annual simulation [15, 16]. In the course of development, it has been shown that further augmentation (cf. II-A4) of the reference time series on the basis of the annual simulation is reasonable, in particular, to increase the number of critical grid states in the dataset in order to achieve a higher accuracy of the trained model. Furthermore, when calculating the optimal power flow, the grid is assumed to be completely observable. For comparison, a state estimation is also calculated in combination with an optimal power flow in order to take account of the incomplete observability and to obtain a reference for the model to be trained here.
@@
Therefore, the training data set consists of tuples of unsolved and solved optimal power flow as well as an extension by state estimation and optimal power flow on reduced observability. The data set generation, therefore, follows a similar approach as presented in [17].
II-A4 Augmentation
To enlarge our training set, we performed data augmentation, where we added noise to the bounds for $P^{\text{min}}$, $P^{\text{max}}$. The set points of each augmented bus are set to its maximum. We augment cases with violations of the lower voltage band to increase their representation to ensure a more balanced distribution of violations in the dataset. We keep the non-augmented initial dataset as test data and train only on the augmented data.
II-B Reinforcement Learning Methodology
A reinforcement learning approach has been developed for end-to-end learning of grid operation decisions, which takes into account the incomplete availability of measurement data in electricity grids. We call this approach grid decision learning (GDL). The methodology for training the reinforcement learning (RL) agent is shown in Fig. 1.
Figure 1: Overview of Model Methodology
The environment has access to all physical properties of the grid. It is predefined which nodes have controllable assets and which do not. The logic works as presented in the following.
Algorithm 1 Reinforcement Learning Training Procedure
1:Define loads and generation for each node from the data set for each time step.
2:Use the Newton-Raphson solver to obtain the current grid state, simulating real-world physics.
3:Generate an observation from the valid grid state, including $P$, $Q$, and $V$ at controllable nodes.
4:RL agent receives the observation and determines the action vector.
5:Apply the action to the environment; recompute grid state using Newton-Raphson.
6:Calculate the reward based on voltage band violations, line overloading, and curtailment costs.
7:Move to the next time step, maintaining the curtailment setting.
8:Pass the reward and the observation back to the agent.
The architecture is designed to adequately learn the curative curtailment decisions on a single grid topology. This is done by resolving grid violations by curtailing active $P$ and reactive power $Q$ at certain buses with flexibility, which are explicitly listed as controllable. Furthermore, the agent is only provided with measurements $P,Q$ and $V$ of observable buses. We assume that every controllable bus is also observable. Additionally, we supply valid ranges $[P_{i}^{\text{min}},P_{i}^{\text{max}}]$ and $[Q_{i}^{\text{min}},Q_{i}^{\text{max}}]$ for each controllable bus $i$. The reinforcement learning agent shall learn the optimal setpoint for $P$ and $Q$ at the controllable buses based on the restricted measurements (or observation). The agent that determines the $P$ and $Q$ setpoints is a neural network obtained from an Actor-Critic model, which we train using Deep Deterministic Policy Gradient (DDPG) [18]. Both Actor and Critic are simple multi-layer perceptrons with two hidden layers, each of width 512. We have implemented the model in PyTorch [19] with TorchRL [20].
II-B1 Environment
@@
The environment can be interpreted as a simulation of the physical world. It is provided with a fixed grid topology from the data set generation process, which gets updated in every training step. The agent interacts with the environment by collecting observations at all observable buses. Based on this observation, the agent learns to apply an action by determining relative $P$ and $Q$ setpoints, which are provided to the environment and update the setpoints in the grid topology. A power flow is computed with a Newton-Raphson procedure to simulate the effect of the agent’s decision. Based on the overall resulting state of the environment, which includes measurements of buses not observable by the agent, a reward is given to the agent.
II-B2 Reward
@@
The reward evaluates the agent’s action on the basis of complete information about the grid status. Let $n$ be the number of buses, $k$ the number of controllable generators and $\ell$ be the number of lines. Let $\mathbf{V}$ be the vector of voltages of the buses in the current state, and denote by $\mathbf{V}^{\text{min}}$ and $\mathbf{V}^{\text{max}}$ the vector of lower and upper limits for the voltage band. We define the voltage loss $\mathcal{L}_{V}\coloneqq\max_{i\in\{1,\ldots,n\}}|\mathbf{V}-\text{clamp}(\mathbf{V},\mathbf{V}^{\text{min}},\mathbf{V}^{\text{max}})|_{i}$ as the maximum deviation from the voltage band. Formally, for vectors $\mathbf{x},\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ we define $\text{clamp}(\mathbf{x},\mathbf{a},\mathbf{b})\coloneqq\max(\mathbf{a},\min(\mathbf{x},\mathbf{b}))$ where $\min$ and $\max$ are applied pointwise. Similarly, for the vector of relative loads $\mathbf{I}$ we let $\mathcal{L}_{I}\coloneqq\max_{i\in\{1,\ldots,\ell\}}\max(\mathbf{I}-\mathbf{1},\mathbf{0})_{i}$ be the maximum relative line load violation. Further, we define the curtailment cost as $\mathcal{C}_{P}\coloneqq\frac{1}{|\mathbf{C}|}\sum_{i=1}^{|\mathbf{C}|}\mathbf{C}_{i}$ where $\mathbf{C}_{i}$ is the amount of power curtailed at the $i$-th controllable generator. The agent obtains a negative reward if the power flow (computed after applying the agent’s actions) does not converge. Otherwise, the agent obtains
@@
II-C Training Procedure
Experiences are collected in a Replay Buffer of size 200,000 from which we sample 100 individual experiences uniformly at random at a time, which are then used to train the Actor and Critic networks.
III Experiments
In this section, we present the experimental results of our method on an example grid from reality. First, we analyse the dataset in III-A and describe our training process in III-B before evaluating the results in III-C.
III-A Data Set Generation
We conduct our experiments on a real low voltage grid of Schleswig-Holstein Netz AG, to which we apply our data generation method described in Section II-A with synthetic load and generation profiles. We defined 7% observable and controllable nodes in the grid as depicted in Figure 2.
Figure 2: Low Voltage grid with 7% observability (provided by Schleswig Holstein Netz AG)
Table I presents the distribution of violations in the generated data set (size: 20,453 quarter-hour supply tasks) and the ratio of resolved violations. Note that violations in either category are non-exclusive.
TABLE I: Distribution of violations in experiment dataset